# MoMuSE: Momentum Multi-modal Target Speaker Extraction for Real-time Scenarios with Impaired Visual Cues

### v1

@@ 4.1 Dataset
4.2 Implementation details
The initial learning rate was set to $1e^{-3}$ when optimizing from scratch and $1e^{-4}$ when initialized from a pre-trained checkpoint, using the Adam optimizer. The learning rate is halved if the best validation loss (BVL) does not improve within six consecutive epochs, and the training stops if the BVL does not improve within ten consecutive epochs. The maximum number of training epochs is set to 100. We set $\theta$ to a relatively large value 0.7 to make sure $\mathbf{e}^{r}_{a}$ is replaced only when $\mathbf{e}^{r}_{c}$ is accurate enough. For online inference, we set $L_{\text{init}}$, $L_{\text{win}}$ and $L_{\text{shift}}$ to 1s, 2.7s and 0.2s, respectively, according to [23]. The hyperparameters $\alpha$, $\beta$, $\gamma$, and $\lambda$ are set to 0.3, 0.7, 0.05, and 0.1, respectively.
The SI-SNR loss is scale-invariant, meaning the energy of the estimated speech may vary across different window steps. To ensure consistent energy levels in the estimated speech, we apply the NeuroHeed normalization strategy [23] in inference.
5 Results
5.1 Causal vs. non-causal
Table 1: SI-SNR (dB) of MuSE with different training settings.
@@ 5 Results
In this section, we explore the optimal training settings for online scenarios where visual cues are impaired. Table 1 demonstrates that the non-causal model configuration [32] outperforms the causal variant in both online and offline evaluation modes. Additionally, incorporating impaired video data into the training set significantly enhances performance in scenarios where visual cues are unreliable. However, this improvement comes at the cost of reduced performance when the visual inputs are normal. This trade-off likely occurs because the model learns to rely less on visual information when it is exposed to impaired visual data during training. The second row of Table 1, which employs a non-causal model and incorporates impaired data during training, shows the best performance in our target scenario, online inference with impaired visual cues. Therefore, this configuration is adopted in subsequent experiments to ensure optimal results.
5.2 Comparative analysis with baselines
Table 2: ‘V’ and ‘I’ denote visual sequences and static image, respectively. ‘Data’ denotes the number of utterances in training set.
@@ 5.2 Comparative analysis with baselines
We evaluate our proposed MoMuSE model and several baseline approaches using multiple performance metrics: Scale-Invariant Signal-to-Noise Ratio (SI-SNR), Signal-to-Distortion Ratio (SDR) [35], Perceptual Evaluation of Speech Quality (PESQ) [36], and Short-Time Objective Intelligibility (STOI) [37]. The results are summarized in Table 2.
@@
5.3 Ablation studies
Fig. 3: Model performance under varied impaired ratios is shown, where MoMuSE_w/o(PI+$\mathcal{L}_{\text{Pe}}$) excludes parameter initialization (PI) and penalty loss ($\mathcal{L}_{\text{Pe}}$), and MoMuSE_w/o($\mathcal{L}_{\text{Pe}}$) excludes only $\mathcal{L}_{\text{Pe}}$.
@@ 5.3 Ablation studies
Table 3: Comparative analysis of various types of impairments.
@@
5.4 Impact of the impairment types
Table 3 presents the performance of MoMuSE and MuSE on three types of visual impairments. In scenarios of complete visual absence, where no useful visual information is available, MoMuSE shows clear effectiveness, highlighting its ability to excel when visual cues are severely compromised. This result underscores the model’s strength in leveraging momentum-based speaker embeddings to maintain robust performance. For cases of lip concealment and low resolution, although MoMuSE does not exhibit significant improvements, its performance remains competitive, suggesting that even with partial visual impairments, the model can effectively utilize available cues while relying on speaker momentum for stability.
5.5 Analyzing total visual absence
Table 4: Comparative analysis on total visual absence. The performance of models on different utterance lengths.
@@ 5.5 Analyzing total visual absence
Table 4 presents the SI-SNR (dB) performance of models with varying utterance lengths. First, MoMuSE consistently outperforms MuSE, demonstrating its ability to maintain focus on the target speaker over time in the absence of visual cues. Second, as the length of the utterance increases, MoMuSE shows improved performance. This improvement can be attributed to its updating mechanism, which discards poorer speaker embeddings and retains more reliable ones in the memory bank during inference. Using these better speaker embeddings, MoMuSE is able to achieve enhanced performance.
An interesting observation is MoMuSE’s performance in Table 4 exceeds its performance in Table 3. We attribute this to the clean visual input during the initialization step in Table 4, whereas impaired visuals might occur in the initialization step in Table 3. This highlights the critical importance of having clean visuals during the initialization step.
6 Conclusion and Future Work

### v2

@@ IV-A Dataset
IV-B Implementation details
The initial learning rate was set to $1e{-3}$ when optimizing from scratch and $1e^{-4}$ when initialized from a pre-trained checkpoint, using the Adam optimizer. The learning rate was halved if the best validation loss (BVL) did not improve within six consecutive epochs, and the training stopped if the BVL did not improve within ten consecutive epochs. The maximum number of training epochs was set to 100. We set $\theta$ to a relatively large value 0.7 to make sure $\mathbf{e}^{r}_{a}$ was replaced only when $\mathbf{e}^{r}_{c}$ was accurate enough. For online inference, we set $L_{\text{init}}$, $L_{\text{win}}$ and $L_{\text{shift}}$ to 1s, 2.7s and 0.2s, respectively, according to [20]. The hyperparameters $\alpha$, $\beta$, $\gamma$, and $\lambda$ were set to 0.3, 0.7, 0.05, and 0.1, respectively.
The SI-SNR loss is scale-invariant, meaning the energy of the estimated speech varies across different window steps. To ensure consistent energy levels in the speech, we applied the NeuroHeed normalization strategy [20] during inference.
V Results
V-A Causal vs. Non-causal
In this section, we explore the optimal training settings for online scenarios where visual cues are impaired.
TABLE I: SI-SNR (dB) of MuSE with different training settings
@@ V Results
Table I demonstrates that the non-causal model configuration [27] outperforms the causal variant in both online and offline evaluation modes. Additionally, incorporating impaired video data into the training set significantly enhances performance in scenarios where visual cues are unreliable. However, this improvement comes at the cost of reduced performance when the visual inputs are normal. This trade-off likely occurs because the model learns to rely less on visual information when it is exposed to impaired visual data during training. The second row of Table I, which employs a non-causal model and incorporates impaired data during training, shows the best performance in our target scenario, online inference with impaired visual cues. Therefore, this configuration is adopted in subsequent experiments to ensure optimal results.
V-B Comparative Analysis with Baselines
We evaluate our proposed MoMuSE model and several baseline approaches using multiple performance metrics: Scale-Invariant Signal-to-Noise Ratio (SI-SNR), Signal-to-Distortion Ratio (SDR) [29], Perceptual Evaluation of Speech Quality (PESQ) [30], and Short-Time Objective Intelligibility (STOI) [31]. The results are summarized in Table II.
TABLE II: ‘V’ and ‘I’ denote visual sequences and static image, respectively. ‘Data’ denotes the number of utterances in training set
@@ V-B Comparative Analysis with Baselines
V-C Ablation Studies
Figure 3: Model performance under varied impaired ratios is shown, where MoMuSE_w/o(PI+$\mathcal{L}_{\text{Pe}}$) excludes parameter initialization (PI) and penalty loss ($\mathcal{L}_{\text{Pe}}$), and MoMuSE_w/o($\mathcal{L}_{\text{Pe}}$) excludes only $\mathcal{L}_{\text{Pe}}$
@@ V-C Ablation Studies
TABLE III: Comparative analysis of various types of impairments
@@
V-D Impact of the Impairment Types
Table III presents the performance of MoMuSE and MuSE on three types of visual impairments. In scenarios of complete visual absence, where no useful visual information is available, MoMuSE shows clear effectiveness, highlighting its ability to excel when visual cues are severely compromised. This result underscores the model’s strength in leveraging momentum-based speaker embeddings to maintain robust performance. For cases of lip concealment and low resolution, although MoMuSE does not exhibit significant improvements, its performance remains competitive, suggesting that even with partial visual impairments, the model can effectively utilize available cues while relying on speaker momentum for stability.
V-E Analyzing Total Visual Absence
TABLE IV: Comparative analysis on total visual absence. The performance of models on different utterance lengths.
@@ V-E Analyzing Total Visual Absence
Table IV presents the SI-SNR (dB) performance of models with varying utterance lengths. First, MoMuSE consistently outperforms MuSE, demonstrating its ability to maintain focus on the target speaker over time in the absence of visual cues. Second, as the length of the utterance increases, MoMuSE shows improved performance. This improvement can be attributed to its updating mechanism, which discards poorer speaker embeddings and retains more reliable ones in the memory bank during inference. Using these better speaker embeddings, MoMuSE is able to achieve enhanced performance.
An interesting observation is MoMuSE’s performance in Table IV exceeds its performance in Table III. We attribute this to the clean visual input during the initialization step in Table IV, whereas impaired visuals might occur in the initialization step in Table III. This highlights the critical importance of having clean visuals during the initialization step.
VI Conclusion and Future Work